OpenStudy (nuccioreggie):

If y − 4 = 2x, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(2, 1), (4, 2), (6, 3)} {(1, 2), (2, 4), (3, 6)} {(4, 0), (6, 1), (8, 2)} {(0, 4), (1, 6), (2, 8)}

OpenStudy (anonymous):

plug the answers in and see which ne you get I just got fussed at for giving direct answers hint hint A

OpenStudy (anonymous):

@suppa_lova i got fussed at too last night

OpenStudy (anonymous):

... but the answer is in the last post as a hint

OpenStudy (texaschic101):

it is not A take a look at D... y - 4 = 2x.....using (0,4)...x = 0 and y = 4 4 - 4 = 2(0) 0 = 0 (correct) y - 4 = 2x...using (1,6)...x = 1 and y = 6 6 - 4 = 2(1) 2 = 2 (correct) y - 4 = 2x...using (2,8)...x = 2 and y = 8 8 - 4 = 2(2) 4 = 4 (correct)